OpenStudy (bananas1234):

Consider the numbers 4/13 and 3/10 Step a compare the numbers. Step b find a rational number between the two given numbers. I also need to show all pf my work.

OpenStudy (bananas1234):

Of*

OpenStudy (bananas1234):

@Ahsome

OpenStudy (bananas1234):

@rational

OpenStudy (anonymous):

13 × 10 10 × 13

OpenStudy (anonymous):

4 x 10 10 10 x 3

OpenStudy (bananas1234):

Sorry, but all that is coming up is a bunch of question marks

OpenStudy (bananas1234):

Ok i can see you second coment

OpenStudy (anonymous):

find a common number for the bottom and that would be 130

OpenStudy (anonymous):

4 /13 = 40 /130 3/10 = 39/130

OpenStudy (anonymous):

39:40 is this helping

OpenStudy (bananas1234):

Yes

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

really me to but this is kinda easy what about this didn't you get?

OpenStudy (anonymous):

at first it I was kinda complicated but if you look at it l Ike multiplication it makes way more sense

OpenStudy (bananas1234):

Not sure if i understand the meaning of a rational number

OpenStudy (anonymous):

ok so a ratio is just a way to compare numbers a:b

OpenStudy (bananas1234):

Ok

OpenStudy (bananas1234):

I am sitll kinda confused

OpenStudy (skullpatrol):